OpenStudy (rocklionmba):

Simplify sine theta over square root of the quantity 1 minus sine squared theta.

OpenStudy (freckles):

recall 1-sin^2(x)=cos^2(x)

OpenStudy (rocklionmba):

you having the problem that you need to leave to write again?

OpenStudy (freckles):

no

OpenStudy (freckles):

are you?

OpenStudy (rocklionmba):

yes its annoying

OpenStudy (freckles):

hmm... not sure why you are having that problem

OpenStudy (freckles):

but do you see that you can use the identity 1-sin^2(x)=cos^2(x)

OpenStudy (rocklionmba):

yes; so the sqare root of 1-sin^2(x) becomes square root of cos^2(x)

OpenStudy (freckles):

\[\frac{\sin(x)}{\sqrt{1-\sin^2(x)}}=\frac{\sin(x)}{\sqrt{\cos^2(x)}}\] now if cos(x) is positive then sqrt(cos^2(x))=cos(x) or if cos(x) is negative then sqrt(cos^2(x))=-cos(x)

OpenStudy (freckles):

\[\text{ so you could say } =\frac{\sin(x)}{\pm \cos(x)} \]

OpenStudy (freckles):

now sin(x)/cos(x)= to what?

OpenStudy (rocklionmba):

exacly because you don't know whats x

OpenStudy (rocklionmba):

wait that's...

OpenStudy (rocklionmba):

nope lost it. it only says to simplify and I think we did

OpenStudy (freckles):

sin(x)/cos(x)=tan(x)

OpenStudy (rocklionmba):

I knew it was something like SOH CAH TOA

OpenStudy (freckles):

=tan(x) if cos(x) is positive =-tan(x) is cos(x) is negative

OpenStudy (rocklionmba):

its both though...

OpenStudy (freckles):

it isn't both it is or

OpenStudy (freckles):

it just depends on what x is

OpenStudy (freckles):

cos(x) is positive in the 1st and 4th quadrant so if x is in the 1st or 4th quadrant then you have tan(x) cos(x) is negative in the 2nd and 3rd quadrant so if x is in the 2nd or 3rd quadrant then you have -tan(x)

OpenStudy (rocklionmba):

it still both negative and positive, but after figuring out the quadrants for sin, would it be negative? im taking a stab at it here

OpenStudy (freckles):

we don't do to consider where sin is pos or neg here

OpenStudy (freckles):

need to *

OpenStudy (rocklionmba):

so then still confused

OpenStudy (freckles):

we had sqrt(cos^2(x)) not sqrt(sin^2(x))

OpenStudy (freckles):

\[\frac{\sin(x)}{\sqrt{1-\sin^2(x)}}=\frac{\sin(x)}{\sqrt{\cos^2(x)}}=\frac{\sin(x)}{\cos(x)} \text{ if } \cos(x)>0 \\ \frac{\sin(x)}{\sqrt{1-\sin^2(x)}}=\frac{\sin(x)}{\sqrt{\cos^2(x)}}=\frac{\sin(x)}{-\cos(x)} \text{ if } \cos(x)<0\] then we can say since sin(x)/cos(x)=tan(x) we have: \[\frac{\sin(x)}{\sqrt{1-\sin^2(x)}}=\frac{\sin(x)}{\sqrt{\cos^2(x)}}=\frac{\sin(x)}{\cos(x)} =\tan(x) \text{ if } \cos(x)>0 \\ \frac{\sin(x)}{\sqrt{1-\sin^2(x)}}=\frac{\sin(x)}{\sqrt{\cos^2(x)}}=\frac{\sin(x)}{-\cos(x)}=-\tan(x) \text{ if } \cos(x)<0\] We can also do one more case... if cos(x)=0 then our fraction we started with doesn't exist you know since we can't divide by zero

OpenStudy (freckles):

so you have three cases: \[\frac{\sin(x)}{\sqrt{1-\sin^2(x)}} =\tan(x) \text{ when } \cos(x)>0 \\ \frac{\sin(x)}{\sqrt{1-\sin^2(x)}}=-\tan(x) \text{ when } \cos(x)<0 \\ \frac{\sin(x)}{\sqrt{1-\sin^2(x)}} \text{ doesn't exist when } \cos(x)=0\]

OpenStudy (rocklionmba):

but we cant say which are true because we don't know cos(x). so then how do we figure it out?

OpenStudy (freckles):

as i said you have the three cases above

OpenStudy (freckles):

all three cases must be mentioned

OpenStudy (rocklionmba):

you know what, I think we went to far. my answers are tan(theta), tan^2(theta),1, and -1.

OpenStudy (freckles):

based on your answers they assumed cos was positive

OpenStudy (rocklionmba):

so then if its positive then your first case is true?

OpenStudy (freckles):

yes since cos is positive

OpenStudy (freckles):

and we know cos is pos because only pos tan was listed

OpenStudy (freckles):

gotta go

OpenStudy (rocklionmba):

ok. thx so much

OpenStudy (rocklionmba):

and you were right too.